Ralph leaves home at 8:00 A.M. for an important business meeting. Twelve minutes later, Mrs. Ralph discovers that Ralph had forgotten his briefcase. Ralph always drives fifty miles per hour. If Mrs. Ralph drives sixty miles per hour, what time will she catch up to Ralph?
Answer by richwmiller(17219)

12 minutes is 1/5 of an hour.
In 12 minutes Ralph goes 10 miles at 50 miles an hour
Since Mrs Ralph goes 10 miles (60-50) an hour faster than Ralph, it will take her an hour to catch up.
r*t=d
Ralph's time is t+1/5
Mrs Ralphs time is t
(t+1/5)*50=t*60
50t+10=60t
10=10t
t=1